sam bought a 6 month $1400 certificate of deposit. At the end of 6 months she received $112 simple interest. What annual rate of interest did the certificate pay?
a 14%
b 8 %
c 13.3 %
d 16%

Answer by stanbon(75887)

Sam bought a 6 month $1400 certificate of deposit. At the end of 6 months she received $112 simple interest. What annual rate of interest did the certificate pay?
--------
I = P*r*t
112 = 1400*r*(1/2)
r = 224/1400
r = 0.16
Ans: 16%
==================
Cheers,
Stan H.
==================
a 14%
b 8 %
c 13.3 %
d 16%